OpenStudy (anonymous):

Seth works 12 hours a week. He recently received a 5% raise. His new weekly pay is $115.92. What was Seth’s weekly pay before the raise? Use decimal multiplier please! What decimal multiplier do I use?

OpenStudy (anonymous):

say that seth's salary before hand was y y * 1.05(decimal multiplier) = 115.92 so y = 115.92/ 1.05 y= 110.4 The reason the multiplier is 1.05 is because you want a five percent increase to the whole value. The whole value is 1 in decimal form and 5% is 0.05 in decimal terms and so the decimal multiplier is 1.05.